The concentrated sulfuric acid we use in the laboratory is 98.0 percent
H2SO4 by mass. Calculate the molality and molarity of the
acid solution. The density of the solution is 1.83 g/mL.
Answer #1

Consider 1 L = 1000 mL of solution

Mass of solution = volume x density

= 1000 x 1.83 = 1830 g

Mass of H2SO4 = 98.0/100 x 1830 = 1793.4 g

Moles of H2SO4 = mass/molar mass of H2SO4

= 1793.4/98.08 = 18.3 mol

Molarity = moles of H2SO4/volume of solution in L

= 18.3/1 = 18.3 M

Mass of water = 1830 - 1793.4 = 36.6 g = 0.0366 kg

Molality = moles of H2SO4/mass of water in kg

= 18.3/0.0366

= 500 mol/kg = 500m

answered by: londan